# LazyDrag: Enabling Stable Drag-Based Editing on Multi-Modal Diffusion Transformers via Explicit Correspondence

### v2

@@ Abstract.
The reliance on implicit point matching via attention has become a core bottleneck in drag-based editing, resulting in a fundamental compromise on weakened inversion strength and costly test-time optimization (TTO). This compromise severely limits the generative capabilities of diffusion models, suppressing high-fidelity inpainting and text-guided creation. In this paper, we introduce LazyDrag, the first drag-based image editing method for Multi-Modal Diffusion Transformers, which directly eliminates the reliance on implicit point matching. In concrete terms, our method generates an explicit correspondence map from user drag inputs as a reliable reference to boost the attention control. This reliable reference opens the potential for a stable full-strength inversion process, which is the first in the drag-based editing task. It obviates the necessity for TTO and unlocks the generative capability of models. Therefore, LazyDrag naturally unifies precise geometric control with text guidance, enabling complex edits that were previously out of reach: opening the mouth of a dog and inpainting its interior, generating new objects like a “tennis ball”, or for ambiguous drags, making context-aware changes like moving a hand into a pocket. Additionally, LazyDrag supports multi-round workflows with simultaneous move and scale operations. Evaluated on the DragBench, our method outperforms baselines in drag accuracy and perceptual quality, as validated by VIEScore and human evaluation. LazyDrag not only establishes new state-of-the-art performance, but also paves a new way to editing paradigms. Here is the project website.
copyright: nonecopyright: none
[…]
1. Introduction
@@
Drag-based editing in diffusion models remains fundamentally challenging. To preserve object identity during editing, prior methods often perform implicit point matching via attention. A common strategy, introduced by MasaCtrl (Cao et al., 2023), shares key and value tokens during attention. However, this strategy allocates more attention weights to spatially nearby regions instead of semantically related ones (Wang et al., 2025a; Feng et al., 2025), which leads to unstable and degrading edits. Rather than tackling this fundamental cause, as a compromise, many methods rely on test-time optimization (TTO) or weakened inversion strength. These compromises mask the mismatch and incur costs, including unreliable inpainting, suppressed text guidance, and distorted edits.
@@
Instead of the compromise, we take a principled alternative: replace implicit attention-based matching with an explicit correspondence map and inject it directly into the generation process. With this reliable map, editing under full-strength inversion becomes stable without TTO, enabling faithful inpainting and text-guided generation. Beyond addressing the fundamental issue, the choice of network architecture remains crucial for editing. The recent transition from U-Nets (Rombach et al., 2022) to Multi-Modal Diffusion Transformers (MM-DiT) (Esser et al., 2024) provides an ideal foundation for this shift, because MM-DiTs offer tighter vision–text fusion, which improves inversion robustness and raises the ceiling for attention control. As shown by ColorCtrl (Yin et al., 2025), this architecture supports stronger semantic consistency and controllability, allowing attention control to be applied across all single-stream attention (SS-Attn) layers without manual selection of specific layer indexes like that in U-Nets. We exploit these advantages by building our method on MM-DiTs.
@@
Unlike in U-Nets, identity preservation in MM-DiTs is non-trivial. Simply sharing key and value tokens, as in DiTCtrl (Cai et al., 2025), does not reproduce the identity-preserving behavior achieved by MasaCtrl with U-Nets (Cao et al., 2023). Recently, CharaConsist (Wang et al., 2025a) showed that re-encoding and injecting semantically aligned tokens can preserve identity in MM-DiTs. However, its point matching relies on the average of attention similarity, which is fragile under full-strength inversion and often yields unsuitable edits. In contrast, drag instructions naturally define a field that maps handle points to target points, forming a deterministic correspondence map. We turn this explicit map into attention controls. This explicit correspondence–driven preservation resolves the root issue, stabilizes edits under full-strength inversion without TTO. As a result, it enhances inpainting and text guidance ability, delivering higher fidelity and controllability than prior methods.
@@
2. Related Work
Figure 2. Effect of inversion strength. Examples of LazyDrag under different inversion strengths. The additional prompt is “a red apple in the mouth”.
@@ 2. Related Work
GAN-based models (Reed et al., 2016; Yu et al., 2023; Wang et al., 2023) have been largely replaced by diffusion models with U-Net backbones (Ho et al., 2020; Rombach et al., 2022) due to better fidelity and stability. However, U-Nets scale poorly, prompting a shift toward Diffusion Transformers (DiT) (Peebles and Xie, 2023). Among them, MM-DiT (Esser et al., 2024) has become the backbone of choice in recent state-of-the-art systems (Esser et al., 2024; AI, 2024; Labs, 2024; Yang et al., 2024; Kong et al., 2024; Liu et al., 2025b), including FLUX (Labs, 2025). We are the first to introduce a drag-based editing method within MM-DiTs.
@@
Training-free text-guided editing methods use pre-trained diffusion models without fine-tuning, offering strong flexibility. Prompt-to-Prompt (Hertz et al., 2023) edits attention maps for localized control, with extensions to images and videos (Wang et al., 2025b; Liu et al., 2024b; Cao et al., 2023; Rout et al., 2025; Xu et al., 2025; Ju et al., 2024). Recent work explores attention control in MM-DiTs: DiTCtrl (Cai et al., 2025) for long video generation, ColorCtrl (Yin et al., 2025) for light-consistent color edits, and CharaConsist (Wang et al., 2025a) for preserving character identity. Modern approaches such as Step1X-Edit (Liu et al., 2025a) and GPT-4o (OpenAI, 2025) have gained popularity due to their efficiency. However, all rely solely on text, which limits spatial precision. We instead introduce a more intuitive and controllable drag-based method.
@@
Drag-based editing enables users to specify explicit spatial transformations by defining source and target points. Existing methods can be divided into two categories: those requiring test-time optimization (TTO), and those that do not. Most prior works fall into the former, beginning with DragGAN (Pan et al., 2023), and expanding to diffusion-based approaches (Shi et al., 2024; Mou et al., 2024a, b; Liu et al., 2024a; Hou et al., 2024; Shin et al., 2024; Zhou et al., 2025; Ling et al., 2024; Zhang et al., 2025). RegionDrag (Lu et al., 2024) extends the interface to support region-level editing. Some methods (Jiang et al., 2025; Choi et al., 2025) incorporate textual prompts to improve semantic understanding, but still suffers from complex instructions. FastDrag (Zhao et al., 2024) is one of only two notable TTO-free methods, achieving faster inference but still falling short of the quality delivered by TTO-based methods. Inpaint4Drag (Lu and Han, 2025) is the other TTO-free method that build on an inpainting model rather than generative model with inversion. However, directly pasting a warped image to fill the edited region introduces strong unnatural warping artifacts. Also, its strong sensitivity to the input mask leads to frequent boundary artifacts and blurring, even with assistance from modern mask generators (e.g., SAM (Kirillov et al., 2023)). Therefore, we adopt a widely used generative model approach with inversion, rather than an inpainting formulation. Additionally, all prior approaches with inversion rely on low inversion strength, which degrades inpainting quality and limits semantic generation. In contrast, we introduce the first drag-based method for MM-DiTs that leverages full-strength inversion and text-guided attention mechanisms, achieving SOTA performance without any per-image tuning.
3. Method
@@ 3.1. Preliminaries
3.1.1. Training-Free Drag Editing in U-Nets: FastDrag.
FastDrag (Zhao et al., 2024) is the first training-free method for drag-based editing, with U-Net models. It has two parts: (1) it computes a displacement field from drag instructions to create an initial latent $\hat{\bm{z}}_{T}$, filling exposed regions via interpolation, and (2) it applies a MasaCtrl-like (Cao et al., 2023) key and value token replacement during self-attention to preserve object identity. However, beyond the implicit locality bias of self-attention, a central trade-off arises: we want handle points to reach their targets while surrounding regions inpaint naturally. Yet after latent initialization, the cue specific to handles is lost, and all moved points are treated uniformly. Forcing exact positional accuracy yields warp artifacts, whereas enforcing naturalness reduces positioning accuracy. Thus, editing accuracy and visual fidelity are in inherent tension. Moreover, its fusion of multiple instructions is brittle: when drags are antagonistic (for example, opening a mouth by moving the upper lip upward and the lower lip downward), averaging the displacements cancels motion near the seam and the mouth fails to open. Moreover, the interpolation used to fill newly exposed regions further replicates nearby textures, producing repeated artifacts in large uncovered areas, as in Fig. 11.
3.1.2. Identity Preservation in MM-DiTs: CharaConsist.
In parallel, CharaConsist (Wang et al., 2025a) introduces identity preservation in MM-DiTs, though it is not an editing method. To enforce identity preservation, it controls attention by concatenating corresponding source tokens into the key (re-encoded) and value tokens and by blending attention outputs. However, its point matching mechanism is critically flawed: it relies on attention similarity to identify matching points between images, a process that is computationally expensive (requiring additional denoising steps) and inherently unstable. Under full-strength inversion, even minor mismatches in the correspondence map can lead to significant visual artifacts, as proved in Tab. 3.
3.1.3. LazyDrag: Bridging the Gap.
LazyDrag bridges the gap between these two methods. We extend the training-free approach of FastDrag from U-Nets to MM-DiTs, while incorporating the fine-grained attention control method of CharaConsist. We resolve the core weaknesses of both approaches by introducing a unified solution: an explicit correspondence map derived from drag instructions. This map provides stable, precise attention controls throughout the generation process, enabling high-quality, accurate edits while avoiding the pitfalls of attention similarity-based matching and the trade-offs inherent in FastDrag.
3.2. Generating the Explicit Correspondence Map
@@
(1) $\begin{gathered}\alpha_{j}^{i}=\begin{cases}\lVert\bm{p}_{j}-\bm{s}_{i}\rVert_{2}^{-1},&\bm{p}_{j}\neq\bm{s}_{i},\\ \infty,&\text{otherwise},\end{cases}\\ \bm{v}_{j}=\bm{v}_{j}^{i^{\star}},\quad\alpha_{j}=\alpha_{j}^{i^{\star}},\quad\text{where}\quad i^{\star}=\operatorname*{arg\,max}_{i}\alpha_{j}^{i}.\end{gathered}$
@@
With the displacement field $\mathcal{V}$ established, we construct the initial latent $\hat{\bm{z}}_{T}$. This process defines our explicit deterministic correspondence map $(\mathcal{M},\mathcal{A})$ and partitions the latent grid into distinct regions for targeted control. First, we define the set of discrete destination coordinates $\mathcal{P}^{\star}=\{\Pi(\bm{p}_{j}+\bm{v}_{j})\mid\bm{p}_{j}\in\mathcal{P}\}$, where $\Pi(\cdot)$ projects to the grid. By resolving collisions where multiple source points map to a single destination $\bm{x}\in\mathcal{P}^{\star}$ (using winner-takes-all), we get the winner index $j^{\star}(\bm{x})=\operatorname*{arg\,max}_{\ j:\ \Pi(\bm{p}_{j}+\bm{v}_{j})=\bm{x}}\ \alpha_{j}$ and formalize our correspondence map: Matching point map, $\mathcal{M}(\bm{x})=\bm{p}_{j^{\star}(\bm{x})}$. Matching weight map, $\mathcal{A}(\bm{x})=\operatorname{min}(1,\alpha_{j^{\star}(\bm{x})})$. Next, we partition the latent space $\Omega$ into four disjoint sets based on the geometry of the warp. These sets correspond directly to the colored regions in Fig. 3 (a): Background $\mathcal{R}^{\mathrm{bg}}$ (gray) that must remain unchanged, Destinations $\mathcal{R}^{\mathrm{dst}}$ (red and blue, a.k.a., $\mathcal{P}^{\star}$ ) where moved content is rendered with identity preserved, Inpainting $\mathcal{R}^{\mathrm{inp}}$ (yellow) initialized from noise, and Transition $\mathcal{R}^{\mathrm{trans}}$ (green) that blends boundaries smoothly. With these regions clearly defined, the updated latent $\hat{\bm{z}}_{T}$ is constructed by applying a specific rule to each region:
(2) $\hat{\bm{z}}_{T}(\bm{x})=\begin{cases}\bm{z}_{T}(\mathcal{M}(\bm{x})),&\text{if }\bm{x}\in\mathcal{R}^{\text{dst}},\\ \bm{\epsilon}(\bm{x}),&\text{if }\bm{x}\in\mathcal{R}^{\text{inp}},\\ \bm{z}_{T}(\bm{x}),&\text{if }\bm{x}\in\mathcal{R}^{\mathrm{bg}}\cup\mathcal{R}^{\mathrm{trans}},\end{cases}$
@@
3.3. Correspondence-Driven Preservation
Having established the explicit correspondence map, we now detail a two-part mechanism operating at the input (Sec. 3.3.1) and output (Sec. 3.3.2) of the attention calculation in single-stream attention layers only (Yin et al., 2025; Deng et al., 2025). Using this map, the mechanism provides fine-grained control that preserves identity and background, ensuring robust full-strength inversion.
3.3.1. Attention Input Control via Token Replacement and Concatenation

### v3

@@ Abstract
The reliance on implicit point matching via attention has become a core bottleneck in drag-based editing, resulting in a fundamental compromise on weakened inversion strength and costly test-time optimization (TTO). This compromise severely limits the generative capabilities, suppressing high-fidelity inpainting and text-guided creation. In this paper, we introduce LazyDrag, the first drag-based image editing method for Multi-Modal Diffusion Transformers, which directly eliminates the reliance on implicit point matching. In concrete terms, our method generates an explicit correspondence map from user drag inputs as a reliable reference to boost the attention control. This reliable reference opens the potential for a stable full-strength inversion process, which is the first in the drag-based editing task. It obviates the necessity for TTO and unlocks the generative capability of models. Therefore, LazyDrag naturally unifies precise geometric control with text guidance, enabling complex edits that were previously out of reach: opening the mouth of a dog and inpainting its interior, generating new objects like a “tennis ball”, or for ambiguous drags, making context-aware changes like moving hands into pockets. Moreover, LazyDrag supports multi-round edits with simultaneous move and scale operations. Evaluated on DragBench, our method outperforms baselines in drag accuracy and perceptual quality, as validated by mean distances, VIEScore and user studies. LazyDrag not only sets new state-of-the-art performance, but also paves a new way to editing paradigms. Here is the project website.
1 Introduction
@@
Drag-based editing in diffusion models remains fundamentally challenging. To preserve object identity during editing, prior methods often perform implicit point matching via attention. A common strategy, introduced by MasaCtrl (Cao et al., 2023), shares key and value tokens during attention. However, this strategy allocates more attention weights to spatially nearby regions instead of semantically related ones (Wang et al., 2025b; Feng et al., 2025), which leads to unstable and degrading edits. Rather than tackling this fundamental cause, as a compromise, many methods rely on test-time optimization (TTO) or weakened inversion strength. These compromises mask the mismatch and incur costs, including unreliable inpainting, suppressed text guidance, and distorted edits.
@@
Instead of the compromise, we take a principled alternative: replace implicit attention-based matching with an explicit correspondence map and inject it directly into the generation process. With this reliable map, editing under full-strength inversion becomes stable without TTO, enabling faithful inpainting and text-guided generation. Beyond addressing the fundamental issue, the choice of network architecture remains crucial for editing. The recent transition from U-Nets (Rombach et al., 2022) to Multi-Modal Diffusion Transformers (MM-DiT) (Esser et al., 2024) provides an ideal foundation for this shift, because MM-DiTs offer tighter vision–text fusion, which improves inversion robustness and raises the ceiling for attention control. As shown by ColorCtrl (Yin et al., 2025b), this architecture supports stronger semantic consistency and controllability, allowing attention control to be applied across all single-stream attention (SS-Attn) layers without manual selection of specific layer indexes like that in U-Nets. We exploit these advantages by building our method on MM-DiTs.
@@
Unlike in U-Nets, identity preservation in MM-DiTs is non-trivial. Simply sharing key and value tokens, as in DiTCtrl (Cai et al., 2025), does not reproduce the identity-preserving behavior achieved by MasaCtrl with U-Nets (Cao et al., 2023). Recently, CharaConsist (Wang et al., 2025b) showed that re-encoding and injecting semantically aligned tokens can preserve identity in MM-DiTs. However, its point matching relies on the average of attention similarity, which is fragile under full-strength inversion and often yields unsuitable edits. In contrast, drag instructions naturally define a field that maps handle points to target points, forming a deterministic correspondence map. We turn this explicit map into attention controls. This explicit correspondence–driven preservation resolves the root issue, stabilizes edits under full-strength inversion without TTO. As a result, it enhances inpainting and text guidance ability, delivering higher fidelity and controllability than prior methods.
Figure 1: (a) Top: Comparison between our method and two baselines. The leftmost image shows the input image with multiple drag instructions, each indicated by a different color. The text below each result indicates the additional prompt used for generation. “N/A” means no additional prompt. TTO denotes test-time optimization, where the method requires fine-tuning per image and multi-step latent optimization per drag instruction. Notably, our method successfully opens the mouth of the dog and inpaints its interior. Furthermore, with prompt guidance, we can generate diverse results even under ambiguous drag inputs without fine-tuning. (b) Bottom: Multi-round editing results using our approach. Our method supports not only sequential drag operations but also simultaneous actions like movement and scaling, maintaining visual coherence throughout.
@@
  • We resolve the ambiguity of drag instructions by coupling the explicit correspondence map with text guidance. This correspondence-driven method preserves identity and background, while enabling natural inpainting and semantically consistent modifications.
@@
2 Related Work
Figure 2: Effect of inversion strength. Examples of LazyDrag under different inversion strengths. The additional prompt is “a red apple in the mouth”.
@@ 2 Related Work
GAN-based models (Reed et al., 2016; Yu et al., 2023; Wang et al., 2023) have been largely replaced by diffusion models with U-Net backbones (Ho et al., 2020; Rombach et al., 2022) due to better fidelity and stability. However, U-Nets scale poorly, prompting a shift toward Diffusion Transformers (DiT) (Peebles and Xie, 2023). Among them, MM-DiT (Esser et al., 2024) has become the backbone of choice in recent state-of-the-art systems (Esser et al., 2024; AI, 2024; Labs, 2024; Yang et al., 2024; Kong et al., 2024; Liu et al., 2025a), including FLUX (Labs, 2025). We are the first to introduce a drag-based editing method within MM-DiTs.
@@
Training-free text-guided editing methods use pre-trained diffusion models without fine-tuning, offering strong flexibility. Prompt-to-Prompt (Hertz et al., 2023) edits attention maps for localized control, with extensions to images and videos (Wang et al., 2025a; Liu et al., 2024b; Cao et al., 2023; Rout et al., 2025; Xu et al., 2025; Ju et al., 2024; Yin et al., 2025a). Recent work explores attention control in MM-DiTs: DiTCtrl (Cai et al., 2025) for long video generation, ColorCtrl (Yin et al., 2025b) for light-consistent color edits, and CharaConsist (Wang et al., 2025b) for preserving character identity. Modern approaches such as Step1X-Edit (Liu et al., 2025b) and GPT-4o (OpenAI, 2025) have gained popularity due to their efficiency. However, all rely solely on text, which limits spatial precision. We instead introduce a more intuitive and controllable drag-based method.
@@
Drag-based editing enables users to specify explicit spatial transformations by defining source and target points. Existing methods can be divided into two categories: those requiring test-time optimization (TTO), and those that do not. Most prior works fall into the former, beginning with DragGAN (Pan et al., 2023), and expanding to diffusion-based approaches (Shi et al., 2024b; Mou et al., 2024a; b; Liu et al., 2024a; Hou et al., 2024; Shin et al., 2024; Zhou et al., 2025; Ling et al., 2024; Zhang et al., 2025; Shi et al., 2024a). RegionDrag (Lu et al., 2024) extends the interface to support region-level editing. Some methods (Jiang et al., 2025; Choi et al., 2025) incorporate textual prompts to improve semantic understanding, but still suffers from complex instructions. FastDrag (Zhao et al., 2024) is one of only two notable TTO-free methods, achieving faster inference but still falling short of the quality delivered by TTO-based methods. Inpaint4Drag (Lu and Han, 2025) is the other TTO-free method that build on an inpainting model rather than generative model with inversion. However, directly pasting a warped image to fill the edited region introduces strong unnatural warping artifacts. Also, its strong sensitivity to the input mask leads to frequent boundary artifacts and blurring, even with assistance from modern mask generators (e.g., SAM (Kirillov et al., 2023)). Therefore, we adopt a widely used generative model approach with inversion, rather than an inpainting formulation. Additionally, all prior approaches with inversion rely on low inversion strength, which degrades inpainting quality and limits semantic generation. In contrast, we introduce the first drag-based method for MM-DiTs that leverages full-strength inversion and text-guided attention mechanisms, achieving SOTA performance without any per-image tuning.
3 Method
@@ 3.1 Preliminaries
3.1.1 Training-Free Drag Editing in U-Nets: FastDrag.
FastDrag (Zhao et al., 2024) is the first training-free method for drag-based editing, with U-Net models. It has two parts: (1) it computes a displacement field from drag instructions to create an initial latent $\hat{\bm{z}}_{T}$, filling exposed regions via interpolation, and (2) it applies a MasaCtrl-like (Cao et al., 2023) key and value token replacement during self-attention to preserve object identity. However, beyond the implicit locality bias of self-attention, a central trade-off arises: we want handle points to reach their targets while surrounding regions inpaint naturally. Yet after latent initialization, the cue specific to handles is lost, and all moved points are treated uniformly. Forcing exact positional accuracy yields warp artifacts, whereas enforcing naturalness reduces positioning accuracy. Thus, editing accuracy and visual fidelity are in inherent tension. Moreover, its fusion of multiple instructions is brittle: when drags are antagonistic (for example, opening a mouth by moving the upper lip upward and the lower lip downward), averaging the displacements cancels motion near the seam and the mouth fails to open. Moreover, the interpolation used to fill newly exposed regions further replicates nearby textures, producing repeated artifacts in large uncovered areas, as shown in Fig. 11.
3.1.2 Identity Preservation in MM-DiTs: CharaConsist.
In parallel, CharaConsist (Wang et al., 2025b) introduces identity preservation in MM-DiTs, though it is not an editing method. To enforce identity preservation, it controls attention by concatenating corresponding source tokens into the key (re-encoded) and value tokens and by blending attention outputs. However, its point matching mechanism is critically flawed: it relies on attention similarity to identify matching points between images, a process that is computationally expensive (requiring additional denoising steps) and inherently unstable. Under full-strength inversion, even minor mismatches in the correspondence map can lead to significant visual artifacts, as proved in Tab. 4.6.
3.1.3 LazyDrag: Bridging the Gap.
Naively extending FastDrag from U-Nets to MM-DiTs and combining it with the attention control methods of CharaConsist exposes and amplifies their respective weaknesses, yielding unusable results, as shown in Fig. 4.6 and Tab. 4.6. LazyDrag resolves these weaknesses with a unified solution: an explicit correspondence map derived from drag instructions. This map provides stable, precise attention control throughout the generation process, enabling high-quality, accurate edits while avoiding the pitfalls of attention-similarity matching and the trade-offs inherent in FastDrag.
3.2 Generating the Explicit Correspondence Map
@@
$\begin{gathered}\alpha_{j}^{i}=\begin{cases}\lVert\bm{p}_{j}-\bm{s}_{i}\rVert_{2}^{-1},&\bm{p}_{j}\neq\bm{s}_{i},\\ \infty,&\text{otherwise},\end{cases}\\ \bm{v}_{j}=\bm{v}_{j}^{i^{\star}},\quad\alpha_{j}=\alpha_{j}^{i^{\star}},\quad\text{where}\quad i^{\star}=\operatorname*{arg\,max}_{i}\alpha_{j}^{i}.\end{gathered}$ (1)
@@
With the displacement field $\mathcal{V}$ established, we construct the initial latent $\hat{\bm{z}}_{T}$. This process defines our explicit deterministic correspondence map $(\mathcal{M},\mathcal{A})$ and partitions the latent grid into distinct regions for targeted control. First, we define the set of discrete destination coordinates $\mathcal{P}^{\star}=\{\Pi(\bm{p}_{j}+\bm{v}_{j})\mid\bm{p}_{j}\in\mathcal{P}\}$, where $\Pi(\cdot)$ projects to the grid. By resolving collisions where multiple source points map to a single destination $\bm{x}\in\mathcal{P}^{\star}$ (using winner-takes-all), we get the winner index $j^{\star}(\bm{x})=\operatorname*{arg\,max}_{\ j:\ \Pi(\bm{p}_{j}+\bm{v}_{j})=\bm{x}}\ \alpha_{j}$ and formalize our correspondence map: Matching point map, $\mathcal{M}(\bm{x})=\bm{p}_{j^{\star}(\bm{x})}$. Matching weight map, $\mathcal{A}(\bm{x})=\operatorname{min}(1,\alpha_{j^{\star}(\bm{x})})$. Next, we partition the latent space $\Omega$ into four disjoint sets based on the geometry of the warp. These sets correspond directly to the colored regions in Fig. 3 (a): Background $\mathcal{R}^{\mathrm{bg}}$ (gray) that must remain unchanged, Destinations $\mathcal{R}^{\mathrm{dst}}$ (red and blue, a.k.a., $\mathcal{P}^{\star}$ ) where moved content is rendered with identity preserved, Inpainting $\mathcal{R}^{\mathrm{inp}}$ (yellow) initialized from noise, and Transition $\mathcal{R}^{\mathrm{trans}}$ (green) that blends boundaries smoothly. With these regions clearly defined, the updated latent $\hat{\bm{z}}_{T}$ is constructed by applying a specific replacement rule to each region:
$\hat{\bm{z}}_{T}(\bm{x})=\begin{cases}\bm{z}_{T}(\mathcal{M}(\bm{x})),&\text{if }\bm{x}\in\mathcal{R}^{\text{dst}},\\ \bm{\epsilon}(\bm{x}),&\text{if }\bm{x}\in\mathcal{R}^{\text{inp}},\\ \bm{z}_{T}(\bm{x}),&\text{if }\bm{x}\in\mathcal{R}^{\mathrm{bg}}\cup\mathcal{R}^{\mathrm{trans}},\end{cases}$ (2)
@@
3.3 Correspondence-Driven Preservation
Having established the explicit correspondence map, we now detail a two-part mechanism operating at the input (Sec. 3.3.1) and output (Sec. 3.3.2) of the attention calculation in single-stream attention layers only (Yin et al., 2025b; Deng et al., 2025). Using this map, the mechanism provides fine-grained control that preserves identity and background, ensuring robust full-strength inversion.
3.3.1 Attention Input Control via Token Replacement and Concatenation